Speed World Cycles sells high-performance motorcycles and motocross racers. One of Speed World's most popular models is the Kazomma 900 dirt bike. During the current year, Speed World purchased eight of these cycles at the following costs:

  

  Purchase Date Units Purchased Unit Cost Total Cost
  July 1 2 $ 49,500 $ 99,000
  July 22 3 50,000 150,000
  Aug. 3 3 51,000 153,000
8 $ 402,000  

   

On July 28, Speed World sold four Kazomma 900 dirt bikes to the Vince Wilson racing team. The remaining four bikes remained in inventory at September 30, the end of Speed World's financial year.

  

Assume that Speed World uses a perpetual inventory system.
a(1)

Compute the cost of goods sold relating to the sale on July 28 and the ending inventory of Kazomma 900 dirt bikes at September 30, using weighted average cost method. (Omit the "$" sign in your response.)

  

  Cost of goods sold $   
  Ending inventory $   

  

a(2)

Compute the cost of goods sold relating to the sale on July 28 and the ending inventory of Kazomma 900 dirt bikes at September 30, using FIFO method. (Omit the "$" sign in your response.)

  

  Cost of goods sold $   
  Ending inventory $   

    

b(1)

Which of the two cost flow assumptions will result in Speed World Cycles reporting the highest profit for the current year?

  

Weighted average cost
LIFO
FIFO

        

b(2)

May Speed World Cycles use the cost flow assumption that results in the highest profit for the current year in its financial statements, but use the cost flow assumption that minimizes taxable income for the current year in its income tax return?

  

Yes
No
Answer #1

PART A1

Cost of Goods Sold = 402000/8*4 = 50250*4 = 201000

Ending Inventory = $ 201000

PART A2

Ending Inventory = 3*51000 + 1*50000 = 203000

COGS = 2*50000 + 2*49500 =  199000

PART B1

FIFO has lower COGS giving higher Net Profit

PART B2

No